OpenStudy (anonymous):

Max is 24 years older than his son, Liam. In two years, Liam will be half as old as Max. How old is Max now? @heather040200 @johnweldon1993

OpenStudy (uri):

@heather040200 o_O stop giving links?

OpenStudy (anonymous):

@uri kkk

OpenStudy (johnweldon1993):

so Max = Liam + 24 M = L + 24 *easier to write* in 2 years...Liam will be 1/2 Max's age so (L + 2) = 1/2(M+2) That is your system of equations....can you solve it?

OpenStudy (uri):

delete it @heather040200

OpenStudy (johnweldon1993):

Oh...well nvm then lol

OpenStudy (anonymous):

what are you talking about??

OpenStudy (anonymous):

we r talking about the link @heather040200